Water flowing along a horizontal pipe has a speed of 2.27 m/s and a pressure of 4.86 atm. Further along, the pipe narrows so that the cross-sectional area decreases by a factor of 6.44. What is the pressure in the narrow section?


1
Expert's answer
2021-01-22T15:05:40-0500

Explanations & Calculations


  • To workout this, it is necessary to assume water flows in laminar flow such that Bernoulli's equation could be applied.
  • Since the pipe is horizontal, the regular Bernoulli equation reduces down to,

"\qquad\qquad\small P_1+\frac{1}{2}\rho v^2_1=P_2+\frac{1}{2}\rho v^2_2"

  • Subjecting the final pressure: "\small P_2" gives,

"\qquad\qquad\small P_2=P_1+\frac{1}{2}\rho(v_1^2-v_2^2)"

  • Then it is about finding the final speed:"\small v_2".


  • The initial pressure is given in atm which needs to be converted into Pascal before substituting,. Then,

"\qquad\qquad\small 4.86atm=4.86\times 1.01325\times 10^5Pa=492439.5Pa"

  • Since the water was assumed incompressible, volumetric flow is consistent thoughout. Then,

"\qquad\qquad\small Q=A_1v_1=A_2v_2"

  • Now "\small A_2=\large\frac{1}{6.44}\small A_1\to\large\frac{A_1}{A_2}=\small6.44"
  • Rearranging for Q gives,

"\qquad\qquad\large\frac{A_1}{A_2}=\large\frac{v_2}{v_1}=\small6.44"

  • Then the final speed is

"\qquad\qquad\small v_2= 6.44\times 2.27ms^{-1}=14.6188ms^{-1}"


  • Therefore, the final pressure is,

"\qquad\qquad
\begin{aligned}
\small P_2&= \small 492439.5Pa+\frac{1}{2}\times 1000kgm^{-3}\times(2.27^2-14.6188^2)\\
&= \small 338461.3Pa\\
&= \small \frac{338461.3Pa}{1.01325\times 10^5}\\
&=\small \bold{3.34atm}
\end{aligned}"